OpenStudy (anonymous):

Solve the equation. 5x + 9 = 1

OpenStudy (katrinakaif):

5x = 1/9 x = 1/9 x 1/5 x = 1/45

OpenStudy (anonymous):

5x + 9 = 1 5x = -9 +1 5x = -8 x = -8/5